assume that a school district has 10,000 6th graders. in the district, the average weight of a 6th grader is 80 pounds, with a standard deviation of 20 pounds. suppose you draw a random sample of 50 students. what is the probability that the average weight of a sampled student will be between 75 and 105 pounds?
Answer by stanbon(75887)

assume that a school district has 10,000 6th graders. in the district, the average weight of a 6th grader is 80 pounds, with a standard deviation of 20 pounds. suppose you draw a random sample of 50 students. what is the probability that the average weight of a sampled student will be between 75 and 105 pounds?
-------
z(75) = (75-80)/[20/sqrt(50)] = -1.7678
-------
z(105) = (105-80)/[20/sqrt(50)] = 8.8388
----
P(75 <= x-bar <= 105) = P(-1.7678 <= z <= 8.8388) = 0.9615
========================
Cheers,
Stan H.